OpenStudy (anonymous):

1/3 = 1/2 Cos^2 (theta), solve for theta

OpenStudy (anonymous):

can someone solve this, showing ALL steps, i am very rusty on my trig functions/algebra

hero (hero):

I will post a response soon. If my browser lets me

OpenStudy (anonymous):

thks!

hero (hero):

\[\frac{1}{3} = \frac{\cos^2(\theta)}{2}\]

hero (hero):

\[3\cos^2(\theta) = 2\]

hero (hero):

\[\cos^2(\theta) = \frac{2}{3}\]

hero (hero):

\[\cos(\theta)^2 = \frac{2}{3}\]

hero (hero):

\[\cos(\theta) = \pm \sqrt{\frac{2}{3}}\]

hero (hero):

\[\theta = \cos^{-1}(\pm \sqrt{\frac{2}{3}})\]

hero (hero):

Does that help?

OpenStudy (anonymous):

yes! you are amazing!

hero (hero):

What values did you get for theta?

hero (hero):

Thanks for the testimonial by the way :)